The following table shows the marginal private benefits and marginal private cost of providing daycare spaces in the province:

Number of Daycare Spaces Marginal Private Benefits ($) Marginal Private Costs ($)
3000 600 300
5000 550 400
7000 500 500
8000 450 600
9000 400 700

Refer to the above information to answer this question. In an effort to increase the number of daycare spaces, suppose the authorities give daycare operators a subsidy of $150 for each child registered. What will be the new equilibrium fee? Select one:
A. $550.

B. $450.

C. $350.

D. $650.

E. $500.
Answer #1

As we know with a subsidy of $150 for each child registered in daycare spaces. This will reduce the Marginal Private costs at each level by $150.

So, Using the above information we can form a new table after reducing $150 as subsidy from Marginal Private costs to get the new Marginal Private Costs.

Number of Daycare Spaces Marginal Private Benefits ($) New Marginal Private Costs ($)
3000 600 300-150 = 150
5000 550 400-150 = 250
7000 500 500-150 = 350
8000 450 600-150 = 450
9000 400 700-150 = 550

As we know equilibrium is attained at the point where,

Marginal Private Benefits = Marginal Private Costs

So, From the above table we can see that Equilibrium is attained when Marginal Private Benefits = Marginal Private Costs = $450.

As, Equilibrium fee is equal to the Marginal Private Benefits or Marginal Private Costs at equilibrium point.

So, Equilibrium fee = $450

  • Hence, Option B. $450 is correct